### v1

# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

Abstract
Small lightweight wave buoys, SFYs, designed to operate near the coast, have been developed. The buoys are designed to record and transmit the full time series of surface acceleration at $52$ Hz. The buoy uses the cellular network to transfer data and position (up to $80$ km from the base station). This reduces costs and increases band-width. The low cost and low weight permits the buoys be deployed easily, and in arrays in areas where satellite and wave models struggle to resolve wave and current interaction. The buoys are tested in a wave-flume, the open water and in the breaking waves of the surf. The conditions range from calm to significant wave heights exceeding $7$ m and crashing breakers with accelerations exceeding $10$ g. The high sample rate captures the impulse of breaking waves, and allow them to be studied in detail. Breaking waves are measured and quantified in the open water. We measure breaking waves in the surf and recover the trajectory of waves breaking in the field to a higher degree than previously done. The time series of surface elevation, and accurate positioning, permits adjacent buoys to be correlated in a coherent phase-resolved way. Finally, we offer an explanation and solution for the ubiquitous low-frequency noise in IMU-based buoys and discuss necessary sampling and design to measure in areas of breaking waves.
1 Introduction
Wave breaking is the primary process by which energy, momentum and gases are exchanged between the air and the sea. It is the limiting factor in wave growth and helps transfer energy through the wave spectrum from shorter to longer waves (Melville, 1996; Babanin, 2009; Deike, 2022). The forces induced by breaking waves are one of the dimensioning factor in marine and nautical engineering, and their effect on coastlines is profound. Most maritime activity happens close to the coast, with more than 60 % of humans living near the coast (Boehm et al., 2017). Estimates further suggest that $80$ of all search and rescue cases are found within $20$ km of the shoreline (Breivik and Allen, 2008). Winds, waves, and currents are heavily modified and influenced by the coastline where waves and currents interact strongly (Halsne et al., 2022, 2023). Consequently, the regions with the most complex wave conditions typically hold the greatest societal importance and are also the most difficult to sample.
Of the many ways to measure waves, in situ heave buoys and wave staffs are the most popular (Holthuijsen, 2010). There exists a range of high–quality commercial moored and drifting wave buoys [see, e.g., the buoys by Datawell (Datawell BV, 2023), SOFAR Ocean (Raghukumar et al., 2019), and Aanderaa (Dorgeville et al., 2018), LainePoiss (Alari et al., 2022)]. These buoys are designed for operational application and therefore have a need to function robustly for long stretches of times. They usually report statistical properties, like significant wave height, and in some cases time series up to a few Hz. Often the time series are only available onboard, and not over telemetry. However, the last decade has seen a small revolution in low-cost, open-source wave-measuring devices (Thomson et al., 2023; Rabault et al., 2020; Feddersen et al., 2023a). These designs generally do not seek to have the same robustness as commercial buoys, but can measure in quite different conditions. Because of the low cost and lower environmental impact, they are also often considered expendable. This is important when measuring in regions where the risk of loss and damage is high. More importantly, these instruments are typically an one or more orders of magnitude smaller and lighter than traditional wave buoys and can thus measure in locations where traditional buoys cannot. Crucially, they also offer the opportunity to measure quantities that larger buoys are incapable of capturing, such as the acceleration in breaking waves (Feddersen et al., 2023b). The limits and capabilities of the instrument determine which questions can be answered. Heterogeneous and different solutions are valuable in themselves in that they have different biases, cover different use cases and make results more robust.
A relevant example to this work is Brown et al. (2018) who previously quantified breakers in the open ocean by measuring accelerations. Similarly, the microSWIFT (Thomson, 2012; Thomson et al., 2019) has been used to study waves near the shore and in areas with strong currents (Rainville, 2022; Rainville et al., 2023). Sinclair (2014) and later Brown and Paasch (2021) and Fisher et al. (2021); Feddersen et al. (2023a) have attempted to measure breaking waves in the surf or in wave laboratories.
In this study we describe and demonstrate a small, lightweight, buoy, dubbed the Small Friendly Buoy (SFY), tailored to coastal waters and designed for capturing and quantifying breaking processes as well as spatially varying wave-current interactions. The code and hardware is open source and available publicly [^1] https://github.com/gauteh/sfy. The work builds on the OpenMetBuoy (Rabault et al., 2020).
Some open-source buoys have telemetry so that the data can be transmitted to shore. This is typically done using Iridium satellites and the Iridium Short Burst Data (SBD) system, which admit only very low bandwidth (except with high-power, heavier, equipment). Some buoy designs can only store the information on board and thus require the buoy to be retrieved. This is often not possible or requires much time and resources. Loss of data is also an issue with such an approach, especially for deployments in rough conditions. The SFY avoids both these problems by relying on the cellular network to transmit data. This allows us to transmit much more data so that we can send continuous high-frequency time series of the full three-dimensional acceleration vector of the buoy. Additionaly, this allows us to build buoys at a much lower cost and with a smaller physical and environmental footprint. Together this makes it possible to analyse the full time series for impulses incurred by breaking waves, and to resolve the phase of waves between buoys deployed in dense spatial networks. In the coastal environment, resolving this spatial variability is as important as (if not more than) having a single high–quality point measurement. The disadvantage is that the deployments must be within the cellular coverage and will not work far from the coast.
Measuring the wave field through acceleration at a point
The working principle of the SFY wave buoy is to measure the acceleration in the wave field, and to convert acceleration as a function of time to displacement as a function of time. This is an established method that has been used for many decades (see e.g. Tucker (1958)), initially with mechanical accelerometers, and more recently with increasingly cheap, accurate and easily available micro-electromechanical system (MEMS) accelerometers and gyroscopes on inertial measurement units (IMU). Since acceleration ($a$) is the second derivative of displacement, we can recover the displacement ($\eta$) by integrating the acceleration twice,
$v(t)=v_{0}+\int_{0}^{t}a(t^{\prime})\mathrm{d}t^{\prime},$ (1a)
$\eta(t)=z_{0}+\int_{0}^{t}v(t^{\prime})\mathrm{d}t^{\prime}.$ (1b)
[…]
$\mathcal{F}\{\eta(t)\}(\omega)=-\omega^{-2}\;\mathcal{F}\{a(t)\}.$ (2)
Here, $\omega=2\pi f$ is the angular frequency [rad/s], $f$ is the linear frequency [Hz] and $\mathcal{F}$ is the Fourier transform. When integrating spectra of acceleration density ($A$) to elevation energy (variance density spectrum), $E$, Eq. (2) becomes
$E(\omega)=\omega^{-4}A(\omega),$ (3)
[…]
$H_{m0}\simeq 4\sqrt{m_{0}}.$ (4)
There are inherent limitations to measuring a single point of the surface, which cannot capture the changing interference patterns produced by the dispersive wave field (Magnusson et al., 1999). These are better resolved with methods that instantaneously span the spatial field, like stereo-video measurements (see, e.g., Benetazzo (2006); Malila et al. (2022); Malila (2022)) or densely sampled point measurements (e.g Donelan et al. (1985, 1996)). Additionally, the motion of wave buoys tend to underestimate the wave crests (Forristall, 2000); moored buoys are affected by the pull of the anchoring and have a tendency to skirt around the highest wave crests, while drifting buoys overestimate the mean water level by spending too much time near the crests.
An inherent assumption of the accelerometer-based approach is that the acceleration experienced by the buoy is the same as the acceleration of the water surface. This assumption starts to break down when the wave oscillations are so rapid that the buoy can no longer reasonably be said to remain at its equilibrium level of immersion. The frequency at which the buoy no longer manages to follow the sea-surface is a function of its shape and buoyancy and can be approximated using the heave response frequency, $f_{\mathrm{hr}}$ (Thomson et al., 2015),
$f_{\mathrm{hr}}$ $=\frac{1}{2\pi}\sqrt{\frac{g\rho V}{ML}}.$ (5)
Here $\rho$ is the density of seawater, $V$ is the displaced volume, $L$ is the height of the buoy, and $M$ is the mass of the buoy and the added mass caused by drag on the water.
The SFY buoy comes in a small free-drifting form and in a larger moored form. The free-drifting buoy has a heave response of approximately $f_{\mathrm{hr}}=3.56$ Hz, while the moored buoy has $f_{\mathrm{hr}}=1.44$ Hz (assuming the buoy is 50 % submerged). Clearly, though, the expression for the heave response is an approximation, and does not take the horizontal length of the buoy into account. At a rule-of-thumb level, one can argue that if the wavelength ($\lambda$) of the wave to be measured is shorter than twice the (horizontal) radius of the buoy, $r$, the buoy will not be able to follow the wave. In deep water, the linear phase speed is $c_{\mathrm{p}}=\sqrt{g/k}$ with $k=2\pi/\lambda$ the wavenumber. From this we find that the geometric cut-off frequency ($f_{\mathrm{gc}}$) where the wavelength equals $2r$ is
$f_{\mathrm{gc}}=\frac{\sqrt{g/k}}{\lambda}=\sqrt{\frac{g}{2\pi\lambda}}.$ (6)
For the free-drifting buoy $f_{\mathrm{gc}}=2.24$ Hz, and for the moored buoy $f_{\mathrm{gc}}=1.09$ Hz. These frequencies will be compared with the spectra from experiments in Sections 3 and 4.
Structure of this paper
This paper is organised as follows. We first present the design of the SFY buoy in Section 2. Wave tank experiments of the buoy response are presented in Section 3. Field experiments from a mooring in open-ocean conditions are presented in Section 4 together with a comparison against a high-resolution wave model. Results from surf-zone deployments are presented in Section 5. Finally, we discuss our findings and compare with other light-weight buoys in Section 6, and present our conclusions in Section 7.
2 Buoy design
The SFY wave buoy consists mainly of three parts; (i) the electronics board with sensors, processing and telemetry (presented in Section 2); (ii) the power source, usually alkaline or lithium batteries (see appendix A); (iii) the housing which protects the other parts and keeps the buoy floating. Various types of housing can be used, but in this work two types are described, an 800 ml bottle for free-drifting buoys, and a $60\times 80$ cm inflatable fender for moored buoys.
Sensor, processing and telemetry
The circular electronics board measures 60 mm in diameter and about 2 cm in height. Figure 1 shows the components with the 800 ml bottle used for housing.
The sensor pipeline generally follows the setup for the OpenMetBuoy (Rabault et al., 2022), hereafter OMB, and benefits from the extensive field testing and validation that the OMB has been subjected to. The inertial measurement unit (IMU) used is either the ISM330DHCX or LSM6DSOX (STMicroelectronics, 2019, 2020), both of which yield gyroscope and accelerometer readings. The IMU low-pass filters the signal internally according to the configured output rate. The theoretical noise level of the ISM330DHCX is typically $60\,\mu g/\mathrm{Hz}^{1/2}$ (maximum $100\,\mu g/\mathrm{Hz}^{1/2}$), according to the manufacturer datasheetSTMicroelectronics (2020). Here, $g=9.81\,\mathrm{m\,s^{-2}}$ is the earth’s gravitational acceleration.
Accelerometer and gyroscope readings from the IMU are fused by applying an Attitude and Heading reference system (AHRS) algorithm implemented with a Kalman filter so that the acceleration is obtained in a North-East-Down (NED) frame of reference (NXP Semiconductors, 2016; Adafruit, 2023). No magnetometer is currently included, so the North-East vector points in an arbitrary direction, but remains stable on short time scales thanks to the measurements provided by the gyroscope. In this analysis only the relative horizontal acceleration is utilized. The IMU is sampled at $208$ Hz and the samples are then fed into the Kalman filter at this same frequency. Finally, the output from the Kalman filter is filtered using a finite-impulse response (FIR) low-pass filter (129 taps with a Hamming window) and decimated down to $52$ Hz or $26$ Hz depending on the application.
It is difficult to quantify the theoretical noise level of the processing chain as a whole since both the accelerometer and gyroscope introduce noise before these are fused in a complex nonlinear way by the Kalman filter. Hence, the noise level of the accelerometer is used as a lower bound of the noise and as an estimate of its frequency dependence. The noise level of the sensors and the processing pipeline as a whole is discussed in more detail by Rabault et al. (2023). Details on how the data are packed, transmitted and received are described in appendix C.
Figure 1: The components that make up the SFY wave buoy. From the left: housing (plastic bottle) with foam for additional stability, battery pack (Alkaline), main board with SD-card, Microcontroller Unit (MCU) in red, modem and GPS in green, IMU, and GPS and cellular antennas.
3 Lab experiment: Wave flume
Experiments were carried out in a wave flume at SINTEF Ocean in Trondheim, Norway (described in Singsaas et al. (2020)). The flume is 12 m long and 0.53 m wide, with a piston-type wave maker from Edinburgh Designs (Edinburgh Designs, 2023) at one end, and a perforated absorbing beach at the other end. The wave maker is programmable and can be set to generate simple harmonic waves as well as wave packets and random realisations of a specified wave spectrum.
The flume is filled with filtered seawater from an intake at approximately 80 m depth in the Trondheim fjord. During the experiments described here, the flume was filled to a depth of 80 cm. Wave gauges can be mounted at any position in the flume, allowing the surface elevation to be sampled at frequencies up to 128 Hz. The wave gauges, also produced by Edinburgh Designs, are based on the principle of measuring resistance between two vertical parallel metal rods. Table 1 describes the configuration of the five buoys and wave gauge used.
Instrument Frequency IMU
IMU Output Real
Wavebug16 52 Hz 52 Hz 48.58 Hz LSM6DSOX
Wavebug23 52 Hz 52 Hz 48.60 Hz LSM6DSOX
Wavebug25 208 Hz 52 Hz 49.07 Hz ISM330DHCX
Wobbler001 52 Hz 52 Hz 55.90 Hz ISM330DHCX
Wobbler002 52 Hz 52 Hz 55.80 Hz ISM330DHCX
Wave gauge 128 Hz 128.00 Hz
Table 1: Buoy and wave gauge configuration for lab experiment.
(a) $f=0.10$ Hz, $A=3.0$ cm (run 6).
(b) $f=0.15$ Hz, $A=5.0$ cm (run 5).
(c) $f=0.25$ Hz, $A=5.0$ cm (run 4).
(d) $f=0.3$ Hz, $A=5.0$ cm (run 3b).
(e) $f=0.5$ Hz, $A=5.0$ cm (run 2).
(f) $f=1$ Hz, $A=5.0$ cm (run 1).
(g) $f=1.5$ Hz, $A=2.0$ cm (run 8a), B16 fixed with rope.
Figure 2: Welch spectrum measured by buoys and wave gauge (WG) measured in a wave flume with piston configured at different frequencies and wave amplitudes. The red vertical line marks frequency where the wavelength is twice the size of the collar width of the buoys ($2.24$ Hz, see Eq. (6)). The vertical dashed lines mark the input frequency and its higher harmonics.
The Welch spectrum in Figure 2 is calculated with a segment length of 4096 for the buoys, and 8192 for the wave gauge which samples at approximately twice the frequency of the wave-buoy.
The measured peaks in the spectra (Figure 2) match the configured piston frequency and its higher harmonics (gray vertical lines) well. The frequency of the wave components measured with the wave gauge also match very well. In the initial experiments the buoys were freely moving at different positions in the wave flume and were subjected to considerable lateral movement and rotation due to Stokes drift, so the phase between the buoys should not be expected to match. Additionally, the vertical amplitude of the generated wave varied considerably laterally along the flume, probably due to interference and reflection (particularly for the lower frequencies).
At approximately $2.2$ Hz there is a marked decrease in the wave-buoy spectra compared to the wave gauge. The collar-width of the buoy (Figure 1) is about $15.5$ cm. From Eq. (6), we see that a wavelength twice that has a deep-water frequency of $2.24$ Hz (red vertical line in Figure 2). The measured amplitude is therefore increasingly dampened as the wavelength approaches the diameter of the buoy. According to Eq. (5), the heave response frequency is $f_{\mathrm{hr}}=3.56$ Hz, but it is very sensitive to the degree of immersion and displacement volume. It seems that the dominating effect on the heave response is the diameter of the buoy, rather than the weight and volume. Yurovsky and Dulov (2020) observe a similar geometric cut-off frequency of about $2.5$ Hz (corresponding to the wavelength twice the hull diameter of their buoy). As they mention, this means that shorter waves can be measured compared to most other alternatives (e.g., $1.2$ Hz for the Spotter buoy (Raghukumar et al., 2019)).
The buoys measure acceleration and have an approximately constant noise level on the measured acceleration. The wave gauge measures displacement and has an approximately constant noise level on the displacement. The difference becomes apparent in the spectrum (Figure 2) at lower frequencies, where the noise level of the wave gauge is flat (horizontal), while the noise level of the wave buoys is proportional to $f^{-2}$ following the double integration from acceleration to displacement. At lower frequencies the wave buoys therefore have a higher noise ratio compared to the wave gauge, whereas at higher frequencies the wave buoy performs better since acceleration is better for measuring high frequency waves (with higher acceleration than displacement), and conversely the wave gauge is better at measuring low-frequency waves. See Sec. 1 and 66.1 for more details.
3.1 Significant wave height
Piston $H_{m0}$ (m)
$f_{p}$ (Hz) Run $A_{p}$ / $H_{m0}$ (m) WG B16 B23 B25 WOB1 WOB2
0.10 6 0.03 / 0.08 0.06 0.01 0.07 0.01 0.01
0.15 5 0.05 / 0.14 0.04 0.17 0.01 0.18 0.18 0.16
0.25 4 0.05 / 0.14 0.11 0.08 0.08 0.09 0.08 0.10
0.30 3a 0.05 / 0.14 0.12 0.13 0.09
0.30 3b 0.05 / 0.14 0.12 0.08 0.08
0.50 2 0.05 / 0.14 0.10 0.12 0.12
1.00 1 0.05 / 0.14 0.15 9.60 0.15
1.50 7a 0.03 / 0.08 0.08 4.68 7.11 4.23 2.14
1.50 7b 0.03 / 0.08 0.08 3.43
1.50 8a 0.02 / 0.06 0.08 1.78
Table 2: Significant wave height measured in the wave flume, automatic high-pass cut-off frequency.
In Table 2 the significant wave height is calculated from the zeroth moment of the elevation variance spectrum ($H_{m0}$, Eq. 4) and compared to the programmed wave amplitude of the piston, and the significant wave height ($H_{m0}$) of the wave gauge (WG). The theoretical $H_{m0}$ for the generated signal with a single harmonic wave with amplitude $A_{p}$ and variance $\frac{1}{2}A_{p}^{2}$ is calculated as (see Holthuijsen (2010), Chapter 3.5):
[…]
$H_{m0}$ $=4\cdot\sqrt{\frac{1}{2}A_{p}^{2}}$ (7b)
The measured signal from the buoys need to be high-pass filtered at a cut-off frequency to avoid low-frequency noise. The wave gauge is not filtered, since it has a flat noise curve on displacement (Figure 2). Automatic selection of low-end cut-off frequency is generally hard. This is a problem that affects IMU based buoys in general (see, e.g., Nose et al. (2018)) and will be discussed further in Section 66.1. Alari et al. (2022) chose the bandpass cut-off frequencies of the measured signal in the wave tank tests to be $0.30$ and $1.28$ Hz and the LainePoiss® buoy was evaluated at frequencies between $0.57$ and $0.73$ Hz.
In Table 2 the cut-off frequency is found automatically by searching for the first spectral minimum above $0.05$ Hz. The algorithm is based on the process outlined in Rabault et al. (2022) (Figure 7) as well as in Tucker (1958); Waseda et al. (2018). However, when this algorithm is unsuccessful it results in unphysically large amplitudes and significant wave height estimates. The auto-detected cut-off frequency is shown and listed in Figure 2. For instance for $1$ Hz in Table 2 and Figure 1(f) the significant wave height for B23 is obviously wrong and overestimated compared to WOB1. This can be attributed to the cut-off frequency being chosen too low for B23, while better for WOB1 (see $f_{c}$ in Figure 1(f), note that the cut-off for B23 is off the scale of the plot). In general it is easier to auto-detect the cut-off frequency when the waves are higher and the signal to noise ratio is greater. Since we know the input frequency for these experiments we can manually choose cut-off frequencies. In Table 3 the significant wave height is shown for the manually chosen frequency band. These values match better, and do not show exaggerated significant wave heights like those present in Table 2 (where a too low or too high cut-off frequency has been picked).
Piston Bandpass $H_{m0}$ (m)
$f_{p}$ (Hz) Run $A_{p}$ / $H_{m0}$ (m) $f_{hp}$ $f_{lp}$ WG B16 B23 B25 WOB1 WOB2
0.10 6 0.03 / 0.08 0.05 1.28 0.06 0.03 0.07 0.03 0.03
0.15 5 0.05 / 0.14 0.05 1.28 0.04 0.17 0.07 0.18 0.18 0.16
0.25 4 0.05 / 0.14 0.10 1.50 0.11 0.08 0.08 0.09 0.08 0.12
0.30 3a 0.05 / 0.14 0.10 1.50 0.12 0.13 0.09
0.30 3b 0.05 / 0.14 0.10 1.50 0.12 0.08 0.08
0.50 2 0.05 / 0.14 0.30 2.00 0.10 0.12 0.12
1.00 1 0.05 / 0.14 0.30 2.00 0.15 0.43 0.15
1.50 7a 0.03 / 0.08 0.30 2.00 0.08 0.18 0.23 0.15 0.08
1.50 7b 0.03 / 0.08 0.30 2.00 0.08 0.11
1.50 8a 0.02 / 0.06 0.30 2.00 0.08 0.10
Table 3: Significant wave height measured in the wave flume, custom bandpass cut-off frequencies.
3.2 Wave buoy orbital measurements
To determine the buoy orientation along the wave-flume we assume that the buoys movement is mainly varying along the wave direction. Since we do not have a magnetometer we do not expect the $x$ and $y$ axes to be pointing in the same direction on long time-scales (limited by gyroscope drift). Figure 3 shows the vertical motion as a function of time for a sine wave of $0.3$ Hz (Figure 1(d)), as well as the vertical and horizontal acceleration.
In order to obtain the trajectory of the wave buoy we apply principal component analysis (PCA) to the high-pass filtered $x$ and $y$ acceleration. The acceleration is filtered to remove drift and only consider the movement due to wave motion. The first principal component, $\bf u_{0}$, is the direction in the $xy$-plane along which the variation is greatest, the second component ($\bf u_{1}$) is orthogonal to $\bf u_{0}$ representing the rest of the variance.
We project the $x$ and $y$-accelerations onto the normalized first principal component ($\frac{\bf u_{0}}{\left\|\bf u_{0}\right\|}$, horizontal) and the normalized second principal component ($\frac{\bf u_{1}}{\left\|\bf u_{1}\right\|}$, vertical), the original movement in the $xy$-plane and the re-projected movement is shown in Figure 4:
[|$\ \bf h_{i}\\ |\end{bmatrix}=\begin{bmatrix}|&|\\ \bf x&\bf y\\ |&|\end{bmatrix}\cdot\frac{\bf u_{i}}{\left\|\bf u_{i}\right\|}$ (8)
$\bf h_{0}$ and $\bf h_{1}$ are the vectors with re-projected $x$ and $y$ accelerations.
If the buoy is only moving along one direction, and the gyroscope is not drifting, all movement should be aligned along the first principal component. It is clearly visible in Figure 4 that this is the case for almost all the movement. The ratio between the two principal components for 10 seconds in the wave-flume is $925$ (upper panel in Figure 4), which provides a measure of how robust this method is on this time scale.
Figure 3: Vertical motion, vertical acceleration and horizontal acceleration along direction of maximum variance for a $0.3$ Hz sine wave. Upper: Vertical displacement for the wave-gauge, B23 and B25. Middle: Vertical acceleration for B23 and B25. Lower: Horizontal acceleration for B23 and B25.
Figure 4: Trajectory of the B25 wave-buoy in the wave-flume with 10 seconds of $0.3$ Hz sinusoidal waves (Figure 1(d)). Upper: xy-plane with principal components and re-projected trajectory. Middle: high-pass filtered trajectory in horizontal–vertical plane along first ($\bf u_{0}$) and second ($\bf u_{1}$) principal component. Lower: full bandwidth trajectory in horizontal–vertical plane for first principal component.
Linear waves predict the orbital motion of particles. If the wave-buoy is to measure the details of wave-motion it should be water-following. The relation between the horizontal and vertical motion, and their spectra, will tell us how well the buoy is following the particle motion in a wave (Feddersen et al., 2023a). The amplitudes of the horizontal and vertical particle orbital velocity in a plane wave are (e.g. Holthuijsen (2010))
$u_{x}$ $=\omega a\frac{\cosh[k(d+z)]}{\sinh(kd)}$ (9)
$u_{z}$ $=\omega a\frac{\sinh[k(d+z)]}{\sinh(kd)}.$ (10)
Here, $d$ the local water depth and $k$ the wave number. The ratio of horizontal to vertical velocity is thus
$\frac{u_{z}}{u_{x}}=\frac{\sinh[k(d+z)]}{\cosh[k(d+z)]}=\tanh[k(d+z)].$ (11)
At the surface, $z=0$, Eq. (11) becomes $\tanh(kd)$. The same relation (11) holds true for the acceleration, $\frac{a_{z}}{a_{x}}$ (differentiate Eq. (9) and Eq. (10) twice). In this setup $k=0.71\;\textrm{m}^{-1}$ for $f=0.3$ Hz, and $\tanh(kd)$ is $0.51$, while the orbital motions in Figure 4 (middle panel) have $\frac{u_{z}}{u_{x}}$ of approximately $0.36$.
3.3 Breaking waves
Figure 5: A breaking wave measured in the wave-flume by the buoys drifting freely next to the wave gauge. The integrated vertical displacement for the wave gauge (WG), and buoys 23 and 25 are shown. The vertical and horizontal acceleration (along main principal component, as in Sec. 3.2). Buoy 23 samples the IMU at $52$ Hz, while buoy 25 samples the IMU at $208$ Hz.
Two observations of the relation between the measured harmonic sine waves and the breaking waves are of particular interest. First, the maximum horizontal acceleration for the sine wave is of the order of ±$\pm 0.05g$ (Figure 3), whereas for the breaking wave in the lab it exceeds ±$\pm 1g$. Second, buoys 23 and 25 show similar accelerations and vertical displacements for the non-breaking wave (upper panel in Figure 3). However, for a breaking wave the integrated vertical displacement (upper panel in Figure 4) is dominated by exaggerated low-frequency signals for buoy 23. The major difference between the two buoys during the recording of the breaking wave is that buoy 23 is sampling the IMU at $52$ Hz, while buoy 25 is sampling at the full $208$ Hz before filtering and down-samplng to $52$ Hz after feeding and re-projecting the acceleration through the AHRS fusion algorithm. These results will be discussed further in Sec. 66.1.
4 Open water: Fedjeosen
A wave buoy built into a fender was deployed at $60.752145^{\circ}$ N, $4.693160^{\circ}$ E, moored at $90$ m depth, west of Hellisøy lighthouse in Fedjeosen in July 2023. The location is exposed to to westerly waves from the North Sea, but is sheltered from the north by Fedje Island. The configuration is described below (shown in Figures 6 and 19) and is very similar to recommended moorings described by Martini et al. (2021).
Fedjeosen is the main shipping lane into Hjeltefjorden and the coastal installations around Bergen. The channel creates a strong tidal current with spatially and temporally varying current gradients across the channel (which is about $2$ km wide). This will give rise to strong wave-current interaction and a spatially inhomogeneous wave field with open-ocean amplitudes. It is thus a challenging location to test the equipment, but also one where both storm waves and low-frequency swell can be found, allowing us to test the full measurement range of the buoy.
The buoy is equipped with $3\times 21=63$ D-cell Alkaline batteries, sufficient for a one-year deployment with transmissions over the cellular network of measurements sampled at $26$ Hz.
(a) Location of mooring in Fedjeosen, map from Kartverket (sjøkart).
(b) A fender of dimensions $65$ cm x $88$ cm floats at the surface, and houses the electronics and batteries.
Figure 6: Fedjeosen mooring.
The buoy was deployed “anchor last” with an anchor of approximately $80$ kg. As significant wave heights of more than $12$ m can be expected during winter storms, more weight is required than in sheltered locations. Still, the weight is almost two orders of magnitude less than for large wave buoys, making the deployment manageable with a smaller vessel. The mooring rope is $12$ mm Danline Polypropylene (PP). This diameter dimension is excessive, and causes unnecessary drag from currents, and a smaller diameter should be chosen in the future. The buoy is deployed by first lowering the floating buoy to the water surface, then the mooring is unspooled while traveling against the weather. Finally the anchor can be lifted off the ship and be dropped with a releasable hook. The rope is weighted with a steel wire of approximately $10$ kg between the surface buoy and a buoyancy float at $30$ m so that the buoy is relatively free to move with the waves, and the rope is kept out of the way for vessels traveling around the buoy. The bottom anchor is split in two parts with some chain as buffer, so that in large waves the first chain and anchor will move first in order to prevent the entire anchor from being lifted and wandering off position (see schematic in appendix D). The cellular modem connects to a station $80$ km away (more details in appendix B).
Figure 7: $H_{m0}$ measured by a moored buoy at the western coast of Norway (Fedjeosen), compared to operational wave models (with and without current) (The Wamdi Group, 1988), satellite observations of significant wave height, and the measured wind nearby.
Figure 8: Observations from the buoy compared to the simulated $H_{m0}$ (WAM800 with current). Linear regression and quantile-quantile (Q-Q) lines are shown. The model predict slightly higher significant wave height, but with a otherwise fairly linear relationship (Pearson correlation coefficient, $R=0.94$). The model without current shows the same $R$-value.
The measured significant wave height is compared to operational coastal wave models (Fig. 7 and 8), with and without currents (The Wamdi Group, 1988). The nearest grid point in the wave model is about $280$ m away from the buoy location. Figure 8 compares the observations against the simulated $H_{m0}$: there is a good match with an $R$-value of $0.94$ for both models. The buoy measures slightly lower $H_{m0}$ values than model predicts, but the relationship is close to linear.
Deploying the SFY at an exposed location like Fedjeosen provides the opportunity to compare its recordings with satellite observations of significant wave height to add an additional reference for the comparison. Measured $H_{m0}$ was compared to the near real time L3 satellite altimeter measurements from the Copernicus Marine Service catalogue product [^2] WAVE_GLO_PHY_SWH_L3_NRT_014_001 (E.U. Copernicus Marine Service Information . Marine Data Store (MDS)(2023), CMEMS). The satellite data are retrieved and collocated with the SFY using the open source tool wavy [^3] https://wavyopen.readthedocs.org following the collocation routine with temporal and spatial constraints as outlined by Bohlinger et al. (2019). To acquire enough data, we set the temporal constraints to ±$\pm 10$ minutes of the time stamp of the significant wave height, and a maximum spatial distance of $100$ km. This resulted in matches with 6 different satellites, namely Sentinel-3A/B, Cryosat-2, CFOSAT, SARAL-AltiKa, and Sentinel-6A Michael Freilich. Satellite observations appear to follow closely the time series of $H_{m0}$ recorded by the SFY (Fig.7).
(a) Hjeltefjorden ($52$ Hz), logarithmic axes.
(b) Fedjeosen ($208$ Hz), logarithmic axes.
(c) Hjeltefjorden ($52$ Hz), linear axes.
(d) Fedjeosen ($208$ Hz), linear axes.
Figure 9: Elevation energy (variance density) spectra for comparable situations (similar $H_{m0}$ and $T_{p}$) of the same buoy (Figure 5(b)) deployed at two different locations. The IMU (before feeding the AHRS-algorithm) was configured to sample at $52$ Hz in the first case (Hjeltefjorden), and at $208$ Hz in the second case (Fedjeosen). The arrow in the lower left panel shows low-frequency noise caused by insufficient sample rate.
In Figure 9 spectra for 20-minute windows of the time series are compared to spectra for the same buoy deployed inside Hjeltefjorden a few months earlier. The deployment was very similar except that for the first deployment the IMU was sampling at $52$ Hz, while for the second (in Fedjeosen) the IMU was sampling at $208$ Hz and downsampled after being passed through the AHRS-algorithm. The spectra (Figure 9) have been chosen for similar wave situation in terms of $H_{m0}$ and $T_{p}$. A notable difference is that for the buoy configured to sample at the lower frequency ($52$ Hz), more low-frequency noise is observed (arrow in Figure 8(a)). This difference will be discussed more in Sec. 66.1.
Between $1$ and $1.1$ Hz, after the $f^{-4}$ slope there is a pole (onset of increased negative logarithmic slope). This corresponds with the frequency of wavelength twice the horizontal diameter of the buoy ($65$ cm) which is $1.09$ Hz. It therefore seems that the dominant factor capping the upper measurable frequency is the size of the buoy.
Breaking waves measured by the moored buoy
Breaking waves hitting or hurling a buoy cause rapid acceleration impulses as can be clearly seen in the wave flume (Figure 5). The strength of this impulse depends on buoy dimensions and wave steepness and how the breaking wave impacts the buoy. Seldal (2023) did a preliminary investigation into the signature of the breaking waves to try distinguish the different types of breakers in a way similar to the approach of Brown and Paasch (2021). The simplest method of quantifying the number of breakers would be to count maxima in the time series of horizontal acceleration, but this would detect many peaks for each breaking event. In order to make the method more robust against false positives we apply the method from Brown et al. (2018). Here we calculate the Short-Time Fourier-Transform (STFT) for tapered windows of $t_{\textrm{bw}}=1.5$ s (overlap between windows of $t_{bw}/2$) for all three components of acceleration ($a_{z}$, $a_{x}$, and $a_{y}$):
$A(f,t^{\prime})=|\mathcal{F}\{a_{z}\}|+|\mathcal{F}\{a_{x}\}|+|\mathcal{F}\{a_{y}\}|\quad\textrm{for time-windows}\;\left<t^{\prime},t^{\prime}+t_{bw}\right>$ (12)
and take the sum ($\alpha$) of the absolute magnitude along the frequency dimension and the sum ($A$) of components for frequencies above $f_{\mathrm{bc}}=3.6$ Hz:
$\alpha(t)=\sum\limits_{f>3.6}A(f,t).$ (13)
Regular, linear waves, do not exist at significant energy levels at frequencies greater than $2$ Hz and we here assume that the energy found at these frequencies is caused by impulsive accelerations from breaking waves. The frequency was chosen so that it is above the consistent band of energy (upper panel in Fig. 9(a)).
(a) Detected breaking events for the moored buoy in Fedjeosen. Upper: The sum of the acceleration spectrum ($A$). Middle: Detected breakers and $\alpha$. Bottom: Breaker fraction ($Q_{b}$) for 20 minute windows.
(b) 2023-10-14T16:00
(c) 2023-10-24T12:00
(d) 2023-11-04T12:15
Figure 10: Breaking events and photos taken towards the buoy position (located to the left behind the hill) at times marked in Figure 9(a). Photos by Norsk Luftambulanse, used with permission.
Breakers are identified by counting peaks of $\alpha$ above a threshold value. The parameters chosen for window length ($t_{\mathrm{bw}}$), $f_{\mathrm{bc}}$ and the threshold for $\alpha$ are dependent on the buoys response to breakers and signal processing (Brown et al., 2018). As in Brown et al. (2018) we use a threshold $\alpha=3.25\,\mathrm{m}\,\mathrm{s}^{-2}$, and rather increase the cut-off frequency to $3.6$ Hz with the argument that only counting high-frequency impulses is a more robust way of detecting the breakers. Along with the plot of $\alpha$ and the detected breakers in the middle panel in Fig. 9(a), we also plot the norm of the acceleration components $\sqrt{a_{z}^{2}+a_{x}^{2}+a_{y}^{2}}$, filtered above $2$ Hz. It gives very similar results to $\alpha$, with the main difference caused by the smoothing due to the windowing in $\alpha$ (middle panel in Fig. 11).
In Figure 11 broadbanded maxima in acceleration (visible as vertical stripes in the acceleration spectrum) are used to detect breakers. In the lower panel the three acceleration components are shown (unfiltered). Similar impulses to the lab experiments are apparent (Fig. 5).
Figure 11: Detected breaking events for the moored buoy in Fedjeosen. Upper: The sum of the acceleration spectrum ($A$). Middle: Detected breakers and $\alpha$. Bottom: Unfiltered acceleration components.
A way to quantify breaking waves is to estimate the breaking frequency (or approximately breaking probability) (Schwendeman and Thomson, 2014; Brown et al., 2018),
$Q_{b}=N_{b}T/\tau.$ (14)
Here, $N_{b}$ is the number of breakers detected during a measurement interval, $\tau=20$ min, and $T$ is the wave period, here approximated with $T_{m01}$. Figure 9(a) shows $Q_{b}$ (lower panel) for a subset of the data. We have visually inspected images taken every 15 minutes with an air traffic camera on the Fedje Island pointing in the direction of the wave buoy. In Figure 10 photos from three instants are shown. These are marked in Figure 9(a) with vertical dashed lines. Qualitatively, the detection of breakers agrees well with the visual impression of breaking waves in the photographs. The wave steepness ($H_{m0}/(g\cdot\frac{T_{m01}^{2}}{2\pi})$) for the measurement period varies very little (but agree well with models).
5 Arrays of wave buoys in the surf
The buoy was tested in the surf in two separate locations for two types of experiments. First to record the waves and breaking events simultaneously on three buoys (Jæren, Norway), and demonstrate that it is possible to study the phase difference of the events between the buoys in the array. Secondly, to try to measure the trajectory of a single buoy as it is caught in a breaking wave (Green Bowl, Bali, Indonesia).
5.1 Array of buoys in the surf
The first three buoys were deployed as a free-drifting array on the beach outside Jæren on the west coast of Norway on the 8th of January 2023. The cross-shore wind and rip currents carried the buoys outwards and they needed to be picked up before drifting too far out at sea. Figure 12 shows the drift trajectory of one of the deployments and the conditions of the day. As the buoys move independently of each other, their relative positions are continuously changing (white dashed triangles).
(a) Drift trajectory with the instantaneous geometry shown as dashed white lines. Aerial images © norgeibilder.no, Kartverket (Rogaland 2019).
(b) Wave conditions on the day. The buoys are visible on the crest, to the left on the wave that is breaking.
Figure 12: Free-drifting array of buoys in the surf outside Jæren.
Figure 13: Vertical acceleration spectrograms for three buoys, captured in the surf at Jæren. Vertical lines of high intensity signals spanning the full frequency spectrum is visible at each buoy, with some delay in phase.
Figure 13 shows the acceleration density spectrogram for the vertical component of the three buoys for drift shown in Figure 12. Vertical lines of high energy impulses assumed to be caused by breaking events are visible, and are correlated in time between the buoys. Since the buoys in this experiment were configured with very sensitive accelerometer range (±$\pm 2g$) and gyroscope range (±$\pm 125$ dps), the sensor saturates and the trajectories are difficult to recover.
5.2 Trajectory in a breaking wave
In a similar experimental setup, a single SFY buoy was deployed in the surf at Green Bowl, Bali, Indonesia, in November 2023. In this deployment the acceleration range is set to ±$\pm 16g$ and the gyroscope range is set to ±$\pm 1000$ dps, and the AHRS filter is run at $208$ Hz. First the spectrum (Figure 14) was measured for 20 minutes outside the breaking waves as a reference. The swell was about 0.9 m while the surfing wave height was about 0.3–0.5 m. The spectrum does not consist of a single peak, so irregular sets of waves breaking onto the shore is expected.
Figure 14: Spectrum captured outside the surf at Green Bowl, Bali.
The buoy was then deployed so that it would drift towards the shore through the breaking waves. The horizontal components are re-projected along the direction of maximum variance as in Sec. 33.2 and integrated to the horizontal and vertical displacement. A cut-off frequency of $0.05$ Hz is used when filtering before integrating.
In Figure 16 a wave approaches the buoy (painted red). The buoy is drawn offshore towards the surf wave (a), and almost escapes the breaking crest (b), but eventually the buoy is caught in the overturning wave (c) and thrown towards the surface of the water (exceeding $10g$ upwards acceleration on impact) and flushed out. A second orbit begins as the next (non-breaking) wave travels towards the shore.
Figure 15: Acceleration and spectrogram at Green Bowl, Bali of a buoy in the breaking waves of the surf.
(a) 22 seconds
(b) 23 seconds
(c) 24 seconds
(d) Orbital movement of the buoy for the time between vertical lines in Figure 15. The black circle marks the start of the buoy trajectory and each subsequent second is marked by a red circle. The black triangle marks the end.
Figure 16: Orbital movement of buoy and photos captured at the same breaking wave recorded by the buoy in Figure 15.
The recorded displacement amplitudes match well with expected values from the spectrum and visual observations. Usually, measurements from breaking waves in the surf are not suitable to be integrated to displacement without heavy filtering. In Figure 15 we can match the photographs with the integrated trace, and see that the recorded trajectory is sensible even with a very low cut-off frequency. This is discussed further in Sec. 66.2.
6 Discussion
6.1 Low-frequency noise and breaking waves
All IMU-based wave buoys will struggle to measure low-frequency signals. The sensor measures acceleration, and the double integration to obtain vertical displacement ((1a) and (1b)) introduces drift which must be removed using detrending. Additionally, a core assumption of AHRS-algorithms is that they measure a high-frequency signal, and very long period waves may not fall into this domain and appear to shift the earth’s gravity vector from the buoys point of view.
The lower frequency signal and noise of the acceleration are amplified compared to the higher frequencies through the integration, Eqs. (2) and (3). Displacement measurements from IMUs are therefore inherently sensitive to low-frequency noise. The low-frequency noise can be seen in Figure 2 and 9, and is similar to spectra measured by other IMU-based buoys (Nose et al., 2023; Chuang et al., 2013; Arraigada and Partl, 2006; Collins et al., 2014; Brown et al., 2018; Veras Guimarães et al., 2018; Rabault et al., 2020; Rainville, 2022). The longest wave periods that can be measured robustly using IMUs are about $20$ seconds ($0.05$ Hz), depending on the conditions (Brown et al., 2018; Van Essen et al., 2018; Datawell BV, 2023). Note that up to $100$ s periods can be observed by using e.g. GPS to measure wave orbital velocities, although this is also not trivial (Thomson et al., 2018). Note that when comparing spectra the length of the signal should be similar in order to expect similar signal to noise ratio.
Consequently, the acceleration must be high-pass filtered with a chosen cut-off frequency. Choosing this cut-off frequency automatically is not trivial, since the assumed low-frequency noise often extends higher than $0.05$ Hz and can be difficult to distinguish from the signal. In this analysis we use a relatively simple algorithm where we attempt to find a minimum in the smoothed spectrum after $0.05$ Hz (Nose et al., 2018; Rabault et al., 2022) which is used as the cut-off frequency (see Sec. 33.1).
In particular, in experiments in the surf-zone we notice that there is often no significant energy minimum between the main peak ($f_{p}$) and the low-frequency noise. The same has been observed for sporadic 20-minute periods in the SFY and OpenMetBuoy when deployed in the open ocean (also observed in other buoys (Nose et al., 2018)). These spectra tend to give exaggerated estimates for $H_{m0}$ or invalid estimates for $T_{p}$ (Alari et al., 2022).
Most other assessments of the performance and behavior of IMU based buoys observe and identify the low-frequency noise (see references above), but its origin is not well explained. Some of it is attributed to the integrated noise of the IMU. However, that does not explain why there is much higher low-frequency noise in some situations than others, or that the noise is well above the expected noise-floor (when extrapolating the noise curve from high to low-frequencies). Alari et al. concluded that the low-frequency energy is misplaced from other frequencies, and that it should not be filtered completely away for calculation of spectral moments as is the usual way to deal with this noise. Brown et al. (2018) also attribute some of the low-frequency noise as alias from higher frequencies. For the MOTUS buoy (Dorgeville et al., 2018), a mechanical filter was used to reduce measured vibrations (from wind, waves, and currents) above the Nyquist-frequency that fold back into lower frequencies. However, for the SFY both the gyroscope and accelerometer have built-in low-pass filters that would avoid sampling alias.
Another concern is that rotation causes centrifugal acceleration. Feddersen et al. (2023a) found that a buoy caught in a overturning breaking wave, with the IMU located fairly close to the center, only undergoes small centrifugal accelerations. They find that angular velocity of a tumbling buoy very seldom exceed $600$ dps during breaking events, and never more than $1000$ dps. The SFY buoy has sampled at ranges of ±$\pm 125$ and ±$\pm 500$ dps in the open-water, and ±$\pm 1000$ dps in the surf zone (a lower range gives a better resolution, unless it saturates).
We observe that the low-frequency noise occurs more frequently in the surf, and only sporadically in the open ocean. As wave breaking is more violent and happens much more often in the surf zone, it seems likely that high-frequency acceleration impulses from breaking waves are the cause of the low-frequency noise in the measured elevation. Specifically, rapid acceleration and rotation caused by a breaking wave insufficiently sampled into the fusion algorithm would cause the buoy to lose orientation and would no longer correctly extract the vertical acceleration. To test this hypothesis we configured buoys with different IMU sample rates ($208$ Hz and $52$ Hz), but equal output rate (filtered using a FIR-filter and decimated). In the wave flume (Sec. 3), both the low-frequency and the high-frequency buoys performed well on harmonic waves (Figure 3). However, when the buoys are subjected to breaking waves in the wave flume (Sec. 33.3), the buoy with the higher IMU-sampling (B25) performs far better than the other (B23), with low-frequency oscillations clearly visible.
A buoy that encounters breaking waves only sporadically yield a spectrum contaminated by these impulses. If this is the case, higher sample rate should better resolve the rotation and acceleration and thus result in less low-frequency noise. The spectra in Fig. 9 suggest that the buoy with $208$ Hz sampling and AHRS frequency performs better on low-frequencies than the one with $52$ Hz sample rate. Detecting and removing segments with breaking impulses from the time series should also improve the spectral estimates.
Additionally, a buoy in breaking waves requires much higher sample rates and ranges in measuring rotation and acceleration. Given the large accelerations measured in breaking waves (several $g$ in the open water, and more than $10g$ in the surf in our case (Sec. 5), and 8148--148 - - 14 reported by Sinclair (2014), Brown et al. (2018), and Feddersen et al. (2023a)), a much higher sample rate is probably required to have a chance at measuring these signals with any success.
In Sec. 5, Fig. 15(d) we were able to retrieve the elevation with a cut-off frequency as low as $0.05$ Hz.
6.2 Capturing and quantifying breaking waves
A breaking wave is a highly impulsive event. Impulses have infinite bandwidth, and it is impossible to fully capture them with a band-limited method. Our experiments in the wave flume (Sec. 33.3), the open water (Sec. 44) and in the surf (Sec. 5) show that the horizontal and vertical acceleration impulse of a breaking wave far exceed those of regular waves. The acceleration experienced by the buoy during a breaking event exceed $10g$, consistent with previous work (Sinclair, 2014; Brown and Paasch, 2021; Feddersen et al., 2023a). An important observation is that the high-acceleration impulses are very short and they will not be adequately captured unless sampled frequently enough.
In the open water (Sec. 4) the breaking fraction is quantified using methods developed by Brown et al. (2018). The fraction qualitatively agrees with photos from the location. However, a more controlled experiment with camera on the buoy, or the buoy located in the field of vision of stereo-video cameras, would provide more reliable context. This would allow better evaluation of how robust this method is and the influence of the parameters that go into the it.
Feddersen et al. (2023a) captures the trajectory of a breaking wave generated at the Surf Ranch remarkably well. In this work we attempt to capture the trajectory in the waves breaking onto a beach (Sec. 5). The waves breaking on the beach do not consist of a single reproducible, relatively clean, wave. Yet, we do capture the trajectory well compared to the spectrum outside the surf and the photographs. The signal was filtered with a cut-off frequency as low as $0.05$ Hz, without resulting in exaggerated displacements. To our knowledge a breaking wave in the ocean has not been captured with a buoy in this way before.
The reason that the buoy can measure this way is presumably due to a combination of using a small wave buoy, high frequency sampling, and a wide measure range on the IMU (accelerometer and gyroscope).
6.3 Cellular network and power source
[…]
7 Conclusions
The SFY, a small, light-weight, drifting or moored wave buoy design with the capability to detect breaking events has been described. The buoy is designed to work in coastal environments and communicates over the cellular network as well as stores the data on SD-card. Its design, code, and hardware is open source and available publicly [^4] https://github.com/gauteh/sfy. The buoy has been proven to work reliably up to $80$ km from the nearest cellular tower, and in its small form measures frequencies from $0.05$ Hz to about $2.2$ Hz (or $1.1$ Hz for the moored buoy, without correction). The main advantages of this buoy are:
  • light-weight ($\approx 0.5$ kg) and low-cost ($\approx 200$ USD): allowing many buoys to be deployed, easy deployment from small vessels or even by swimmers or surfers.
  • high-bandwidth: with a greater bandwidth for data transmission, the full time series of AHRS-processed high-frequency accelerations ($52$ Hz) is transmitted back. This allows a much more detailed on-land analysis compared to transmitting only the wave spectra.
  • open source: anyone can build and modify the buoy to adapt it to their experiment.
Through wave flume experiments, deployments in the open-water, and in the surf, the buoy is shown to measure significant wave height and other spectral properties accurately. It is shown to measure the surface elevation in such detail that the individual wave phases can be compared between adjacent buoys. These properties enable the buoy to be used in array configurations where the spatial details of the wave field should be studied, e.g., within a stereo-video footprint (Malila, 2022) or in areas with strong wave-current interactions.
The most challenging measurements were performed in the surf zone. The irregular waves breaking on the beach were not produced by a singular wave component or soliton. Yet, by measuring at very high frequency ($208$ Hz) and with a large dynamic range (±$\pm 16g$ and ±$\pm 1000$ dps), we were able to reconstruct the trajectories of the buoy through breaking and non-breaking waves on the beach with a higher degree of success than previously done. Necessary to this is that we were able to use a low cut-off frequency of $0.05$ Hz here, normally reserved for open-waters. The lack of exaggerated elevation amplitudes after integrating to displacement gives us confidence that the buoy measures correctly. In Table 4 we recommend parameter ranges depending on the deployment scenario.
Scenario Frequency Acceleration Gyroscope
Sea-ice $100$ Hz ±$\pm 2g$$\pm 125$ dps
Open-water $208$ Hz ±$\pm 4g$$\pm 500$ dps
Surf $>208$ Hz >±$>\pm 16g$$\pm 1000$ dps
Table 4: Recommended sample rate and IMU range for different scenarios (see also Sinclair (2014); Rabault et al. (2020); Feddersen et al. (2023a)).
An important conclusion which we surmise will be valid for all IMU-based buoys is that low-frequency noise can to a large degree be explained by impulsive events, like breakers, that are not adequately sampled. The inadequate sampling causes the AHRS algorithm to become confused, and the large, but very short, acceleration peaks are not well captured. We find that by sampling the IMU and running the AHRS-algorithm at a higher frequency before downsampling its output (if bandwidth is limited), much better noise characteristics are achieved. By this logic, better noise characteristics or estimates based on the spectrum (like $H_{m0}$) may improve if impulsive events can be detected and excluded from the spectrum.
The advent of small, disposable drifting wave buoys with telemetry opens new opportunities for dense sampling of the surf zone and regions with strong spatial variability, such as tidal currents in open-ocean conditions (see, e.g., the studies of the wave modulation by the Moskenes tidal current in Northern Norway by Saetra et al. 2021; Halsne et al. 2022, 2023). Deploying large, densely spaced, arrays of buoys that sample at rates high enough to identify breakers and the accelerations in breaking events should allow for experiments that are currently not possible, or prohibitively impractical and expensive to do.
Acknowledgements.
[…]
Appendix A Power usage
The majority of the power is drawn by the cellular modem during connection and transmission of data. When the modem is connecting, current spikes up to $2$ A can occur for a few milliseconds at a time. Super-capacitors limit this to about $0.3$ A, and thus enable batteries with higher internal resistance to be used. Figure 17 shows the measured current (at $4.2$ V) during cellular connection, transmission, and lower-current logging for a buoy configured to log at $26$ Hz and with storage on the secure digital (SD) card temporarily disabled (as in Section 4). The average power consumption for this configuration is about $32$ mW with an average current going as low as $0.6$ mA when logging to memory only. Note that spikes at the beginning of each section are artefacts caused by measurement range change in measuring the current.
Considerable effort has gone into making the buoy work well with alkaline batteries rather than lithium-based batteries. Lithium batteries can deliver much greater power and capacity, but mixing lithium batteries with water is not safe (Hasvold et al., 2007). Since this buoy is designed to work close to the coast and to be drifting freely there will always be an appreciable chance that the buoy is found and picked up by the public. This puts stricter requirements on power-use and maximum current draw.
We found the buoys lasted about $12$ days with continuous logging and transmission during field experiments with a logging frequency of $52$ Hz, storage to an SD card and frequent GPS fixes. This autonomy was obtained with 6 Alkaline C-cells with individual capacity of 8 Ah @ 1.5V (Duracell, 2019), or 16 Ah @ 4.5 V with $2\times 3$ batteries. That amounts to an average current use of about $\mathrm{16Ah/(12\times 24)h=0.055A=55mA}$. However, when the buoy is configured to take positions every hour, without an SD-card, and with a lower output frequency, the current drain is reduced down to about $7.8$ mA (or $32.8$ mW).
Figure 17: Current measured using a CurrentRanger. The periods with the lowest current happen when the buoy is logging the IMU continuously, while not doing anything else. Higher current is drawn when the GPS or cellular communication is active. Spikes seen at the start of new communication or GPS periods are measurement errors due to input-range change from $\mu$A to mA in the CurrentRanger.
A moored version of the buoy (as described in Section 4) with $3\times 21$ Alkaline D-cells (21 parallel $4.5$ V) shows an initial voltage drop of $2.47$ mV on average each day for the first $60$ days. The voltage drop of Alkaline cells is steeper in the beginning and the end the battery capacity. A conservative estimate of the battery capacity with a working voltage between $4.621$ and $3.650$ V, yields a battery time of about $390$ days.
Appendix B Cellular network range
Figure 18 shows a histogram of the distance to cellular tower used by the moored version of the buoy. Most of the time it uses a tower about $80$ km away.
Figure 18: The distance to the cellular tower from the buoy.
[…]
Appendix C IMU sampling and data collection
Due to chip-to-chip differences every IMU samples at a slightly different frequency (typically within ±$\pm 10\%$). The actual sample rate can be retrieved from the IMU registers (STMicroelectronics, 2021) (section 6.4). However, since the data is accurately timestamped using GPS, we estimate the true frequency accurately by the median time difference between every 1024-sample batch. This is done routinely to ensure correct sampling frequency.
The MCU configures the IMU to record acceleration and rotation into its first-in, first-out (FIFO) queue continuously, and then drains and processes the samples at sufficiently frequent intervals for the FIFO to not fill up. The MCU sleeps in between, saving power.
The data are collected in batches of 1024 samples and transmitted in messages tagged with position and timestamps. These data are relayed to a server that stores the messages in a database. A set of processing scripts read all messages in a given range and concatenate the data and construct a continuous time coordinate. The data are stored in CF-compliant NetCDF or zarr files with all metadata, and should be fully self contained and ready for further analysis.
[…]
Appendix D Fedjeosen mooring
Figure 19: Schematic of mooring configuration in Fedjeosen.